OpenStudy (anonymous):

∫e^(7x)sin(2x)dx , Find the integral?

OpenStudy (loser66):

let u = sin (2x) and dv = e^(7x) dx repeat the process again you will have the answer.

OpenStudy (anonymous):

\[\int\limits_{0.1}^{0.7}cosx \ln(2sinx)dx\] what about this one ?